OpenStudy (anonymous):

Giving Medal and fanning who answeres CORRECTLY! 16 + x - 11= 18 - 5 Solve for x

pooja195 (pooja195):

16+x−11=18−5 16+x+−11=18+−5 (x)+(16+−11)=(18+−5)(Combine Like Terms) x+5=13 x+5=13 subtract 5 from both sides to solve for x

OpenStudy (anonymous):

x= 8.... ty @pooja195